OpenStudy (anonymous):

rewrite the expression with positive exponets. 3x^-5, 8x^-7y^-8, (4x)^-2, 1/(3x)^-3

OpenStudy (anonymous):

\[\left\{\frac{3}{x^5},\frac{8}{x^7 y^8},\frac{1}{16 x^2},27 x^3\right\} \]